OpenStudy (anonymous):

Is the following always, sometimes, or never true? 6+8x-9=11x+14-3x

OpenStudy (mathstudent55):

Add like terms on each side. Can you do that?

OpenStudy (anonymous):

No. Im not good with these type of problems.

OpenStudy (mathstudent55):

Like terms are terms that have the same variable part. On the left side, there are three terms, 6, 8x, and -9. 6 and -9 have the same variable part (no variable at all). That means 6 and -9 can be added together. What is 6 - 9 = ?

OpenStudy (anonymous):

-3

OpenStudy (mathstudent55):

Great. The 6 and -9 of the left side are combined into -3. On the left side we now have 8x - 3 Now let's deal with the right side.

OpenStudy (mathstudent55):

On the right side there are also three terms, 11x, 14, -3x In this case, 11x and -3x are like terms because they have the same variable part (each one has x in it). 11x and -3x can be combined together. What is 11x - 3x = ?

OpenStudy (anonymous):

8x

OpenStudy (mathstudent55):

Good. The right side, then is 8x + 14. Let's put together the two sides in our equation: 8x - 3 = 8x + 14 Now we subtract 8x from both sides: We get -3 = 14 Since this is a false statement, that means the equation is never true. In other words, there is no value of x that will male the statement true.